OpenStudy (anonymous):

how can i solve for x in sqrt of 2x-4 minus sqrt of x-5 = 3

OpenStudy (anonymous):

maybe \[\sqrt{2x-4}-\sqrt{x-5}=2\]

OpenStudy (anonymous):

yeah thats right, but its = 3 not 2

OpenStudy (anonymous):

start with \[\sqrt{2x-4}=\sqrt{x-5}+3\] and square both sides

OpenStudy (anonymous):

this is going to be a drag because you are going to have to square twice, and also check your answer. first time you get \[2x-4=x-5+9+6\sqrt{x-5}\]

OpenStudy (anonymous):

this gives \[x-8=6\sqrt{x-5}\] and then square again to get \[(x-8)^2=36(x-5)\] \[x^2-16x+64=36x-180\] \[x^2-52x+244=0\]

OpenStudy (anonymous):

then i guess you can solve using the quadratic formula, but the numbers are so annoyingly large it might be easier to complete the square. your choice

OpenStudy (anonymous):

thanks